OpenStudy (anonymous):

Write an equation of the line that passes the given point and has the slope slope. Do the inverse of that equation as well. (0,4), m=2

OpenStudy (anonymous):

has the slope slope?

OpenStudy (anonymous):

ooh i see slope is 2, point is \((0,4)\) so you know the \(y\) intercept is 4 \(y=mx+b\) with \(m=2,b=4\)

OpenStudy (anonymous):

i got y-4 = 2 (x-0) but how do you do the inverse?